OpenStudy (anonymous):

Expand each logarithmic expression: ⦁ logb(35/43) ⦁ Logb(3x2) Write the expression as a single logarithm: c) 8logbx – 3logby

OpenStudy (solomonzelman):

\[\huge\color{blue}{ Log_b(\frac{35}{43}) ~~~and~~~ Log_b(3x^2) ~~~?}\]

OpenStudy (solomonzelman):

\[\huge\color{red}{ Identities: }\] \[\huge\color{red}{ Log_x(r/y) =Log_xr -Log_xy}\]\[\huge\color{red}{ Log_x(r \times y) =Log_xr +Log_xy}\]\[\huge\color{red}{ Log_x(r ^ y) =yLog_xr }\]

OpenStudy (triciaal):

when you divide numbers subtract the logs log b (35/43) =log b 35 - log b 43

OpenStudy (triciaal):

log b (3x2) multiply numbers add the logs log b 3 + log b 2

OpenStudy (triciaal):

Write the expression as a single logarithm: c) 8logbx – 3logby the number in front is the power to which the number is raised to subtract log you divide the numbers log b (X^8)/(y^3)

OpenStudy (anonymous):

Triciaal can u help me with my new open questiom